OpenStudy (anonymous):

A store sells batteries in packets of 6 or 10. It has a stock of 25 packets containing a total of 186 batteries. How many of these are 10-battery packets?

OpenStudy (anonymous):

s = 6 pack batteries t = 10 pack batteries s + t = 25 -->s = 25 - t 6s + 10t = 186 now sub 25 - t in for s in the 2nd equation 6(25 - t) + 10t = 186 150 - 6t + 10t = 186 150 + 4t = 186 4t = 186 - 150 4t = 36 t = 9 s + t = 25 s + 9 = 25 s = 25 - 9 s = 16 check... 6s + 10t = 186 6(16) + 10(9) = 186 96 + 90 = 186 186 = 186 (correct) There were 9 ten pack batteries = 90 batteries There were 16 six pack batteries = 96 batteries ------------------------------------------ total : 25 packs of batteries = 186 batteries